Encircle the odd one of the following 
 (a) (–1, –2) (b) (–5, +2) (c) (–4, +1) (d) (–9, +7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here
–1 + (–2) = -3–5 + 2 = -3
–4 + 1 = -3
–9 + 7 = -2
Hence, (–9, +7) is the odd one.

Thanks